wo sinusoidal waves of wavelength λ = 2/3 m and amplitude A = 6 cm and differing with their phase constant, are travelling to the right with same velocity v = 50 m/s. The resultant wave function y_res (x,t) will have the form:


y_res (x,t) = 12(cm) cos⁡(φ/2) sin⁡(150πx-3πt+φ/2).


y_res (x,t) = 12(cm) cos⁡(φ/2) sin⁡(3πx+150πt+φ/2).


y_res (x,t) = 12(cm) cos⁡(φ/2) sin⁡(150πx+3πt+φ/2).


y_res (x,t) = 12(cm) cos⁡(φ/2) sin⁡(3πx-180πt+φ/2).


y_res (x,t) = 12(cm) cos⁡(φ/2) sin⁡(3πx-150πt+φ/2).

Clear selection

Two identical sinusoidal waves with w


1
Expert's answer
2021-05-24T16:48:15-0400

Let two waves, "y_1=A\sin(kx-\omega t)" and "y_2=A\sin(kx-\omega t+\phi_o)" be travelling in the same direction,

"\phi_o" : Phase difference

Resultant wave, "y=y_1+y_2"

"y=A\sin(kx-\omega t)+A\sin(kx-\omega t+\phi_o)"

"y=A\sin(kx\pm\omega t+\alpha)"

"\tan \alpha=\dfrac{\sin\phi_o}{1+\cos\phi_o}"

From question,

"A=6\space cm"

"\lambda=\dfrac{2}{3}\space m"

"v=50\space m/s"

"\omega=\dfrac{2\pi v}{\lambda}=150\pi"

"k=\dfrac{2\pi}{\lambda}=3\pi"

"\phi_o=" Phase constant = "3\pi"

"\tan \alpha=\dfrac{\sin3\pi}{1+\cos3\pi}"

"\alpha=\tan^{-1}\infin"

"\alpha=\dfrac{\pi}{2}"

As wave travels in left direction equation of wave will be

"y=A\sin(kx+\omega t+\alpha)"

"y=6\sin\bigg(3\pi x+150\pi t+\dfrac{\pi}{2}\bigg)"